# Aligning Reasoning LLMs for Materials Discovery with Physics-aware Rejection Sampling

### v1

@@ 4.2 Results
Our experiments show that PaRS effectively optimize the teacher’ reasoning traces to induce reasoning capability for property prediction on student model. We present our findings in two parts: first, an analysis of the trace selection process itself, and second, an evaluation of the trained student LRMs on property prediction task.
@@
We assess rejection sampling strategies on traces generated by the teacher Qwen3-235B. As summarized in Table 1, PaRS achieves lower prediction error while requiring fewer generations on average, placing it on the empirical quality–efficiency Pareto front (Fig. 4). A simple MAE gate could make low error appear trivial, yet an external judge that is not optimized for our metric still assigns PaRS the highest overall score. Self-consistency, which often lowers MAE, receives the weakest judge score among baselines. This gap suggests that generic preference signals do not fully align with physics-grounded proximity to ground truth. By combining physics-aware gates with temperature scheduling and early stopping, PaRS concentrates supervision on high-fidelity, physically admissible traces rather than on merely “good-looking” rationales.
Evaluation of trained LRMs
Training the student Qwen3-32B on traces selected by our method yields consistent gains in accuracy, correlation with ground truth, and physical admissibility, while using substantially fewer supervision traces than competing approaches (See Table 2). The LLM-as-a-Judge baseline is competitive in error and correlation but does not match the reduction in violations or the calibration gains achieved by PaRS. Retaining all traces increases supervision volume yet underperforms, consistent with amplified label and reasoning noise when traces remain unfiltered.
@@
4.3 Discussion
This result clarifies why general LRM training strategies are hard to directly transfer to process aware property prediction. Recent work (muennighoff2025s1, ) shows that curated rationales and test time scaling can yield strong results even when only about 53% of traces are correct, likely because tasks such as math, coding, and QA have binary correctness and explicit derivations. For property prediction, categorical correctness is a weak proxy for supervision quality because two traces may follow different discrete steps yet end within 1% of the true EQE, which is the signal the model must learn. OpenThoughts (guha2025openthoughts, ) finds that multi sampling often helps by preserving reasoning diversity, but for recipe to property the same unfiltered diversity amplifies label and trace noise, increases physics violation rate and weakens calibration. PaRS reconciles these views by keeping diversity where it matters, namely multiple near correct physical pathways, while trimming supervision to numerically consistent traces that respect simple physics. The result is a better accuracy and admissibility trade-off with about $15\times$ fewer traces than the multi sampling baseline.

### v2

@@ 4.2 Results
Our experiments show that PaRS effectively optimize the teacher’ reasoning traces to induce reasoning capability for property prediction in student model. We present our findings in two parts: first, an analysis of the trace selection process itself and second, an evaluation of the trained student LRMs on the property prediction task.
@@
We evaluate rejection sampling strategies on the traces generated by the teacher Qwen3-235B. As summarized in Table 1, PaRS achieves lower prediction error while requiring fewer generations on average, placing it on the empirical quality–efficiency Pareto front (Fig. 4). A simple MAE gate could make low error appear trivial, yet an external LLM-as-a-judge that is not optimized for our metric still assigns PaRS the highest overall score. Self-consistency, which got the lowest MAE except for our method, receives the weakest judge score among the baselines. This gap suggests that generic preference signals do not fully align with physics-grounded proximity to ground truth. By combining physics-aware gates with temperature scheduling and early stopping, PaRS concentrates supervision on high-fidelity, physically admissible traces rather than on merely “good-looking” rationales.
Evaluation of distilled LRMs
Training the student Qwen3-32B on traces selected by our method yields consistent gains in accuracy, correlation with ground truth, and physical admissibility, while using substantially fewer supervision traces than competing approaches (See Table 2). The LLM-as-a-Judge baseline is competitive in error and correlation but does not match the reduction in violations or the calibration gains achieved by PaRS. Retaining all sampled traces enlarges supervision volume yet underperforms, consistent with amplified label and reasoning noise when traces remain unfiltered. Recent work (Guha et al., 2025) finds that multi-sampling can help by preserving reasoning diversity, but in our task the same unfiltered diversity amplifies trace noise, increases physics violation rate and weakens calibration. PaRS reconciles these views by keeping diversity where it matters, namely multiple near correct physical pathways, while trimming supervision to numerically consistent traces that respect simple physics. The result is a better accuracy and admissibility with about $15\times$ fewer traces than the multi sampling baseline.
@@
Table 3: Comparison of reasoning traces generate by three teacher models. We calculate MAE for all traces by using all generated traces (seven traces per prompt in average). MAE from selected traces is computed only over selected traces by PaRS. Selected ratio is the fraction of generated traces accepted by textttPaRS.
Teacher model MAE for all traces MAE for selected traces Selected ratio
DeepSeek-R1-671B $2.144$ $0.785$ $0.789$
Qwen3-235B $2.327$ $0.829$ $0.798$
Qwen3-235B-FP8 $9.078$ $0.812$ $0.811$
Table 4: Distillation results. Each row fine tunes the same student (Qwen3-32B) on selected traces generated by the indicated teacher models with PaRS.
Distiled models # train MAE $\boldsymbol{R^{2}}$ Spearman $\boldsymbol{\rho}$ Viol. (%) DeepSeek-R1-Distill-Qwen-32B $7897$ $1.755$ $0.445$ $0.717$ $25.1$ Qwen3-235B-Distill-Qwen-32B $7980$ $1.808$ $0.424$ $0.705$ $27.7$ Qwen3-235B-FP8-Distill-Qwen-32B $8112$ $1.801$ $0.413$ $0.712$ $28.9$
4.3 Analysis
Robustness of PaRS across teacher models
To validate robustness of our method across the diverse teacher models, we sample reasoning traces with DeepSeek-R1-671B and Qwen3-235B-FP8 in addition to Qwen3-235B. MAE for all traces in Table 3 shows that smaller or low-precision teacher models produce noisier trace distributions than large models. However, after applying PaRS, the selected traces from small quantized teachers converge to the quality of those of large teachers: error and stability tighten, and the retained supervision is physically admissible and consistently high quality. In effect, PaRS equalizes teacher induced variability by filtering out unstable generations and preserving only high fidelity reasoning trajectories.
This equalization propagates to the student. Fine-tuning the same student on selected traces from different teachers yields similar MAE, correlation, and violation rates. The correctness ratio of selected traces also varies little across teachers, indicating that PaRS is broadly compatible with heterogeneous teacher models and precisions. Consequently, even smaller or low-precision teachers can provide supervision competitive with that from larger models.
0.30.30.50.50.70.70.90.91.81.8222.22.22.42.4Training correctness ratio rrMAE ($p$)MAE ($p$)Phys. viol. (%)2626282830303232343436363838Physics violations (%)
Figure 5: Effect of training correctness ratio on student performance. Training dataset size and inference token budget are fixed. Points show means over a 5-model ensemble on test prompts.
How trace correctness ratio shapes distillation performance
We study how the fraction of accepted traces (the “correctness ratio” under the tolerance gate $\varepsilon_{\mathrm{MAE}}$) relates to student performance (Fig. 5). Holding the dataset size and token budget fixed, increasing the acceptance rate $r$ monotonically lowers MAE and reduces physics violations. Interestingly, recent work (Muennighoff et al., 2025) shows that curated rationales and test-time scaling can yield strong results even when only 53% of traces are correct. We hypothesize that tasks such as math, coding, and QA exhibit binary correctness with explicit derivations, so trace correctness closely mirrors supervision quality. However, for property prediction, categorical correctness is a weak proxy for supervision quality because two traces may follow different discrete steps yet end within 1% of the true EQE, which is the signal the model must learn.
Table 5: Ablation on adaptive halting. We use 10 nodes of 8 $\times$ H100 (80G) to sample reasoning traces from the 10k prompts with Qwen3-235B and measure the compute time.
Method $K_{\mathrm{avg}}$ MAE Compute time (h)
PaRS w/o Adaptive Halting $8.9$ $0.817$ $20.5$
PaRS w/ Adaptive Halting $6.4$ $0.829$ $15.1$
Ablation on adaptive halting
We ablate the adaptive-halting mechanism in PaRS to assess the benefit of early stopping. As summarized in Table 5, enabling adaptive halting reduces the average number of sampled candidates from $K_{\mathrm{avg}}{=}8.9$ to $6.4$ and shortens compute time from $20.5$,h to $15.1$,h, with only a marginal change in accuracy. These results indicate that the variance- and improvement-based early stopping criteria are effective, low-overhead heuristics for pruning unproductive samples without materially degrading the quality of the retained traces. As the teacher model’s parameters are fixed, persistently low variance and minimal improvement suggest limited capacity to produce faithful reasoning traces that recover ground-truth properties with physically plausible values.